In a rectangle, one side equals x+10 and its opposite side equals 3y-3. The other side equals 2y and its opposite side equals x+7. Find x and y.
Answer by amarjeeth123(574)

x+10=3y-3
x+13=3y..........equation 1
2y=x+7
x+7=2y...........equation 2
Subtracting equation 2 from equation 1 we get
13-7=3y-2y
y=6.
x+7=12
x=5.